OpenStudy (anonymous):

Help??? the last part confuses me. Select the proper equation in point-slope form for the given point and slope. Point: (–3, 7); Slope: 4 y – 7 = 4(x – 3) y – 7 = 4(x + 3) y + 7 = 4(x + 3) y + 7 = 4x – 12

OpenStudy (yttrium):

What do you think is the answer? Any try.

OpenStudy (anonymous):

Guessing doesn't help me...

OpenStudy (anonymous):

Maybe the easiest way iof doing this problem is to plug in the coordinates. (-3,7) There is only 1 equation that satisfies it.

OpenStudy (yttrium):

Yep. It's the same as the problem you posted earlier. y-y1 = m (x-x1) y1 = 7 x1 = -3 m = 4 y-7 = 4 (x-(-3)) Therefore, what should it be @HarukaN

OpenStudy (anonymous):

y – 7 = 4(x – 3)

OpenStudy (yttrium):

Nope. @HarukaN Take note, it is x-(-3)

OpenStudy (owlcoffee):

Let's recall a formula wich I won't prove because i'm too lazy: \[y-y _{1}=m(x-x _{1})\] where x1 and y1 are the coordinates of the point and m is the slope. we have a point and the slope, all we have to do is plug them in: \[y-7=4(x-(-3))\] and simplifying the (-)(-): \[y-7=4(x+3)\]

OpenStudy (anonymous):

I tried relating this problem to the other one I had with your explaination but i ended up confusing myself @Yttrium

OpenStudy (anonymous):

@Owlcoffee can i borrow your brain please. thanks c: